OpenStudy (anonymous):

describe the domain of g(x,y)= 3x +5y/(x^2 +y^2-9)

OpenStudy (anonymous):

Domain = all POSSIBLE numbers. With that being said, you cannot divide by 0. Therefore whatever x and y values that make the denominator = zero are NOT in the domain. Set x^2+y^2-9 equal to zero and solve. Since there are two variables you must must solve for y first and then subsitute that in for y.

OpenStudy (anonymous):

i got x= (7)^1/2 & y=.125

OpenStudy (anonymous):

I havent dont the math but if those two numbers make the denominator zero. the domain is x cannot equal (7)^1/2 & y=.125

OpenStudy (anonymous):

ok, thanks